# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v2

@@ 5 Conclusions
In this paper, based on Halpern iteration and gradient RED, we propose HPPP and GraRED-HP3 for IR problems, which can converge strongly to a particular fixed point $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$. Numerical experiments verify the regularity of HPPP, and the effectiveness GraRED-HP3 for image deblurring and inpainting, which can achieve better performance than classic algorithms with TV regularization. In the future, we plan to study the convergence of nonconvex implicit regularization [25] and extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [3].
Appendix A Boundedness and $\mathcal{M}$-projection
Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important Lemma from [23, Lemma 2.5].
@@ Appendix A Boundedness and $\mathcal{M}$-projection
Let $\mathcal{T}$ be a $\mathcal{M}$-FNE operator, $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz, and $\{\mathbf{u}^{k}\}$ be the sequence generated by (5), $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$, and $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, which satisfies the following:
@@
    $\{\mathbf{u}^{k}\},\{\mathcal{T}\mathbf{u}^{k}\}(k\in\mathbb{N})$ are bounded;
@@
    $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0,k\to\infty$;
@@
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|\to 0,\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0,k\to\infty$.
@@
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$, $C=L\left\|\mathcal{C}\right\|$. For $\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime}\in\mathcal{H}$,
@@
$\leq L\left\|\mathcal{C}\right\|\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}=C\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}$
For (i), $\mathbf{u}^{k}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$,
@@
$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|,C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}<+\infty.$
and
$\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\leq C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}<+\infty$
So is the sequence $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
For (ii), we should show $\mathcal{T}$ is $\mathcal{M}$-asymptotically regular $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0$.
$\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$. There exists a positive real number $M>0$ such that
$\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq M,$
it follows from the $\mathcal{M}$-FNE operator $\mathcal{T}$ and boundedness of $\{\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\}_{k\in\mathbb{N}}$,
(23) $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
@@
$\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|M$
@@
If $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0$ from Lemma A.1.
(24) $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}+\mathbf{u}^{k+1}-\mathcal{T}u^{k}\right\|_{\mathcal{M}}$
$\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|_{\mathcal{M}}+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}.$
Thus, $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0$.
For (iii), there exists $M^{\prime}$ such that $\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k-1}\right\|\leq M^{\prime}$,
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1}\right\|+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k-1}\right\|$
$\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|,$
Therefore $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|\to 0$.
@@
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|\to 0,k\to\infty.$
Lemma A.4.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$, and $\mathrm{Fix}(\mathcal{T})$ be a closed convex subset of a Hilbert space $\mathcal{H}$, and $\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function. There exists a unique solution $u^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
@@
(25) $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
Proof A.5.
@@
Let $l(\mathbf{u})=\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is the differentiable convex function. Assume that $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ is the optimal solution, $\mathrm{Fix}(\mathcal{T})$ is the convex set, thus $t\in(0,1),\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for $\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$,
@@
$=\left\langle 2\mathcal{M}(\mathbf{u}^{*}-\mathbf{a}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
$=2\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0.$
If $\mathbf{u}^{**}$ is the another solution such that $\left\langle\mathbf{u}^{**}-a,\mathbf{u}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0$. Replace $\mathbf{u}$ with $\mathbf{u}^{**},\mathbf{u}^{*}$ in the above two inequalities, respectively.
(26) $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}^{**}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,$
(27) $\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
If we add two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. Since
$\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$
It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$ from above inequality.
As mentioned in [2, 9] and uniqueness projection onto $\mathrm{Fix}(\mathcal{T})$, we can introduce the following notion of $\mathcal{M}$-projection.
Definition A.6 ($\mathcal{M}$-projection).
Assume $\forall\mathbf{u}^{0}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{u}^{0}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{u}^{0}\right\|_{\mathcal{M}}(\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}))$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}-$projection of $\mathbf{u}^{0}$ onto $\mathrm{Fix}(\mathcal{T})$, also denoted $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{u}^{0})$.
Lemma A.7.
Let $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ such that $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$ and $\mathbf{u}^{0}\in\mathcal{H}$, then the following conditions are equivalent:
@@
    $\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{u}^{0})$;
@@
    $\left\langle\mathbf{u}^{*}-\mathbf{u}^{0},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
Proof A.8.
See Lemma A.4 and Definition A.6.

### v4

@@ 5 Conclusions
This paper introduces a Halpern-type PPP (HPPP) algorithm, which has the advantage of strong convergence over PPP. In addition, GraRED-HP3 is proposed by integrating HPPP with denoiser priors for IR problems, which is an accelerated PnP-ADMM. Advantages of HPPP are demonstrated through several toy examples, and GraRED-HP3 achieves state-of-the-art performance on IR experiments. In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [4].
Acknowledgments
The authors thank Prof. Defeng Sun for bringing the interesting work [52] to our attention and Mr. Feng Xue for his insightful discussions. We sincerely appreciate Associate Editor Michael Elad and the anonymous referees for their thorough review and valuable comments, which have substantially improved this paper.
Appendix A Boundedness
Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important lemma from [30, Lemma 2.5].
@@ Appendix A Boundedness
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by (5). Then the following assertions are hold:
@@
    The sequences $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ are bounded;
@@
    If the sequence $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$;
@@
    If the conditions of (ii) hold, then $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0,\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|=0$.
@@
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [9, Proposition 2.3]. Denote
(25) $C=L\left\|\mathcal{C}\right\|,$
for $\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime}\in\mathcal{H}$, we have
@@
$=\left\|(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}(\mathcal{C}^{*}\mathbf{u}^{\prime}-\mathcal{C}^{*}\mathbf{u}^{\prime\prime})\right\|$
$\leq L\left\|\mathcal{C}\right\|\left\|\mathcal{C}^{*}(\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime})\right\|$
$=L\left\|\mathcal{C}\right\|\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}$
(26) $=C\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}},$
where the third inequality follows from the $L$-Lipschitz property of $(\mathcal{M}+\mathcal{A})^{-1}$, and the fourth equality follows from the definition of $\left\|\mathbf{u}\right\|_{\mathcal{M}}$, i.e., $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}=\left\|\mathcal{C}^{*}\mathbf{u}\right\|$ for any $\mathbf{u}\in\mathcal{H}$.
First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$, we have
@@
Denote
(27) $C_{1}=\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\},$
then we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq C_{1}$ for all $k\in\mathbb{N}$.
@@
$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|,CC_{1}\}<+\infty.$
By (26), we have
(28) $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\leq C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq CC_{1}<+\infty.$
So the sequence $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is also bounded.
For (ii), we will prove $\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$. Let $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}$, by the triangle inequality and the $\mathcal{M}$-non-expansiveness of $\mathcal{T}$, we have
$\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$
$\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
(29) $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}=M.$
Note that $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
@@
(30) $\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
where the last inequality follows from Lemma 2.3 and (29).
@@
Applying Lemma A.1 together with $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$ to (30), we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$. Since $\mathbf{u}^{k+1}-\mathbf{u}^{k}=\mu_{k+1}(\mathbf{a}-\mathcal{T}\mathbf{u}^{k})$, we have
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mathbf{u}^{k}-\mathbf{u}^{k+1})+(\mathbf{u}^{k+1}-\mathcal{T}{\mathbf{u}^{k}})\right\|_{\mathcal{M}}$
(31) $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|_{\mathcal{M}}+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}.$
Using (29) and $\lim_{k\to\infty}\mu_{k}=0$ in (31), we can obtain $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$.
For (iii), let $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, by the triangle inequality and (28), we have
(32) $\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+\left\|\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}=M^{\prime}.$
Similar to the proof of (30), we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1}\right\|+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1}\right\|$
(33) $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|,$
where the last inequality follows from (26). Applying (30) and $\lim_{k\to\infty}\mu_{k}=0$ to the above inequality (33), we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0$. Similar to the proof of (31),
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|.$
@@
By $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0,\lim_{k\to\infty}\mu_{k}=0$ and boundedness of $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$, we finally prove $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=0$.
Appendix B $\mathcal{M}$-projection
Lemma B.1.
@@ Appendix B $\mathcal{M}$-projection
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Then there exists a unique solution ${\mathbf{u}}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
@@
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
Proof B.2.
@@
Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function. Assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$. Since $\mathrm{Fix}(\mathcal{T})=\mathrm{zer}\mathcal{A}$ is the convex set, thus $\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ and $t\in(0,1)$. Hence,
@@
$=\left\langle\mathcal{M}(\mathbf{u}^{*}-\mathbf{a}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
(34) $=\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0.$
If $\mathbf{u}^{**}$ is the another solution, then it also satisfies
(35) $\left\langle\mathbf{u}^{**}-{\mathbf{a}},\mathbf{u}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
Replace $\mathbf{u}$ with $\mathbf{u}^{**},\mathbf{u}^{*}$ in (34) and (35), respectively, then
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}^{**}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,$
$\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
Add above two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
$\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0.$
It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$, which completes the proof of Lemma B.1.
As mentioned in [3, 15], we can introduce the following notion of $\mathcal{M}$-projection.
Definition B.3 ($\mathcal{M}$-projection).
Assume $\mathbf{a}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
The following Lemma B.4 extends the variational inequality of the metric projection.
Lemma B.4.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Then the following assertions are equivalent:
@@
    $\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$;
@@
    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
Proof B.5.
See Lemma B.1 and Definition B.3.